OpenStudy (anonymous):

Can someone please help me! Find a particular solution of the differential equation (-5/4)y''−2y'+y=2xe^(4x) using the Method of Undetermined Coefficients (primes indicate derivatives with respect to x).

OpenStudy (amistre64):

constant coefficients eh

OpenStudy (anonymous):

yea i ended up getting (20/729)e^(4x)-(2/27)xe^(4x) and its wrong!

OpenStudy (anonymous):

I tried it 3 times and got the same answer

OpenStudy (amistre64):

-5/4 r^2 -2r +1 = 0 is the characteristic eq which makes this thing look like it might go ugly

OpenStudy (anonymous):

but we only need to solve for the particular solution

OpenStudy (amistre64):

\[r=\frac{2\pm\sqrt{4-4(-5/4)}}{-5/2}\] \[r=\frac{2\pm\sqrt{4+5}}{-5/2}\] \[r=\frac{2\pm3}{-5/2}\to\ \frac{4\pm 6}{-5}\]

OpenStudy (amistre64):

cant solve for p without c can you?

OpenStudy (anonymous):

I thought you didntreally need to solve for the complimentary solution unless you were solving for the whole solution

OpenStudy (anonymous):

but i may be wrong

OpenStudy (amistre64):

i cant verify that, i always remember doing it this way :)

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

\[\large y_c=c_1exp({-\frac45 x})+c_2exp({\frac65 x})\]

OpenStudy (anonymous):

what I did was that i found yp=Ae^4x +Bxe^4x then i derived that and derived it again then I plugged it into the original equation and then solved for the coefficients A and B

OpenStudy (amistre64):

your way way might be right, its been awhile since i tried to figure out the specifics

OpenStudy (amistre64):

id have to read up on undetermines to be sure tho

OpenStudy (anonymous):

but my answer ended up being wrong so please continue on ith hat you were doing

OpenStudy (amistre64):

well, my way might get to an answer, but its prolly the wrong method its asking for :)

OpenStudy (anonymous):

true but if I know the correct answer then I might be able to figure out where I went wrong, so if you can, please continnue

OpenStudy (amistre64):

\[\large y_p=Aexp({-\frac45 x})+Bexp({\frac65 x})\] \[\large y'_p=\cancel{A'exp({-\frac45 x})+B'exp({\frac65 x})}^0\]\[\ \ \ \ \large -A\frac45 exp({-\frac45 x})+B\frac65exp({\frac65 x})\] \[\large y''_p=-A'\frac45 exp({-\frac45 x})+B'\frac65exp({\frac65 x})\]\[\large \ \ \ \ +A\frac{16}{25} exp({-\frac45 x})+B'\frac{36}{25}exp({\frac65 x})\] then its alot of fractioning math lol

OpenStudy (amistre64):

im outof time for today, but good luck. to double chk your answers, use the wolf. http://wolframalpha.com

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=%28-5%2F4%29y%27%27%E2%88%922y%27%2By%3D2xe%5E%284x%29 looks like you got a -2x/27 e^4x

OpenStudy (anonymous):

thanks